What is the variance of the following given sample of daily travel expenses of audit staff: 130, 50, 90, 80, 70, 125, 180, 80, 110?


1
Expert's answer
2021-12-14T03:14:53-0500

Mean "\mu=\frac{130+50+90+80+70+125+180+80+110}{9}=\frac{905}{9}=100.56."

Sample variance:

"\sigma^2=\frac{1}{9-1}[(130-100.56)^2+(50-100.56)^2+(90-100.56)^2+(80-100.56)^2+(70-100.56)^2+(125-100.56)^2+(180-100.56)^2+(80-100.56)^2+(110-100.56)^2]=1627.78."